The perimeter of a rectangle is 166 m. The length is 8 m more than twice the width. What is the length? What is the width?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

The perimeter of a rectangle is 166 m. The length is 8 m more than twice the width. What is the length? What is the width?
let width be x
length = 2x+8
perimeter = 2*(l+w)
= 2*(x+2x+8)
=2x+4x+16=166
6x=150
x=25m
length = 58m

Answer by stanbon(75887)

The perimeter of a rectangle is 166 m. The length is 8 m more than twice the width. What is the length? What is the width?
------------------
P = 2(W + L)
166 = 2(W + (2W+8))
83 = 3W + 8
3W = 75
W = 25 meters (width)
L = 2W+8 = 58 meters (length)
=================================
Cheers,
Stan H.